# Advancing Complex Wide-Area Scene Understanding with Hierarchical Coresets Selection

### v1

@@ Abstract.
Scene understanding is one of the core tasks in computer vision, aiming to extract semantic information from images to identify objects, scene categories, and their interrelationships. Although advancements in Vision-Language Models (VLMs), have driven progress in this field, existing VLMs still face challenges in adaptation to unseen complex wide-area scenes. To address the challenges, this paper proposes a Hierarchical Coresets Selection (HCS) mechanism to advance the adaptation of VLMs in complex wide-area scene understanding. It progressively refines the selected regions based on the proposed theoretically guaranteed importance function, which considers utility, representativeness, robustness, and synergy. Without requiring additional fine-tuning, HCS enables VLMs to achieve rapid understandings of unseen scenes at any scale using minimal interpretable regions while mitigating insufficient feature density. HCS is a plug-and-play method that is compatible with any VLM. Experiments demonstrate that HCS achieves superior performance and universality in various tasks. The code is available at https://wangjingyao07.github.io/HCS.github.io/
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06submissionid: 204ccs: Computing methodologies Scene understandingccs: Computing methodologies Transfer learningccs: Computing methodologies Feature selection
@@ 1. Introduction
Despite their robust representational capabilities (Fu et al., 2024; Cao et al., 2024; Wang et al., 2023b), existing VLMs struggle with wide-area scenes. These scenes—spanning expansive environments such as deep-sea regions, remote sensing imagery, and areas with complex geographical features (Greene and Oliva, 2009; Gu et al., 2019)—differ substantially from the urban or indoor settings typical of VLM pre-training (Naseer et al., 2018; Grant and Flynn, 2017). The challenges primarily arise from two interrelated factors. First, the extensive coverage inherent to wide-area scenes introduces an unprecedented diversity of object types, including novel organisms (e.g., unknown species of sea snakes, ascidians, or sea urchins in oceans) that are rarely encountered in standard datasets. Second, although the overall species variety is high, the data may exhibit a pronounced long-tail and sparse distribution in some areas. For instance, while sedimentary rocks might dominate the seabed, rarer features such as sponges appear only sporadically, leading to an overrepresentation of homogenized, high-frequency features. One might argue that existing VLMs boast an abundance of parameters and deeper network architectures that facilitate the capture of intricate and global semantic information. However, these advantages do not inherently overcome the data challenges present in wide-area scenes. In fact, the models may overly focus on globally uniform, high-frequency features. Recent studies have proposed sample partitioning and multi-scale attention mechanisms (Shi et al., 2024; Wang, 2024; Zhang et al., 2024a; Wang et al., 2024c) to address these shortcomings; however, their effectiveness is heavily dependent on the quality of regional segmentation and scale priors, and they often neglect the relationships between different regions. Furthermore, the substantial computational resources required to process the massive volumes of visual data further restrict the practical deployment of such models. Therefore, while pre-training endows VLMs with impressive capabilities, the unique challenges of wide-area scenes pose significant obstacles to their effective adaptation.
@@
To tackle this, we aim to propose a method that can adaptively refine the relationships between scene regions and model decisions in a coarse-to-fine manner for complex wide-area scene understanding. This allows the VLMs to quickly grasp scenes of any scale or complexity using only a few interrelated and interpretable regions, thus achieving fast understanding without specific training. We begin by investigating the interpretable regional characteristics that support effective selection in wide-area scenes. Inspired by data compression theory (Storer, 1987; Sayood, 2017), we propose that the essence of feature selection is to find a small, weighted subset of a large dataset, ensuring that learning tasks on this subset yields an approximately optimal solution to the original dataset. Accordingly, accurate scene understanding requires that the model establish appropriate evaluation criteria tailored to specific tasks, such as object classification or semantic segmentation, and filter out regions that critically influence decision-making to achieve the key subset, i.e., interpretable regions. To achieve this, we incorporate coreset theory (Har-Peled and Mazumdar, 2004), which enables the construction of a compact subset of input data, i.e., coreset, that serves as a proxy for the whole dataset while maintaining minimal deviation in accuracy. By reducing the data volume, coresets enhance scalability and efficiency. A common approach for constructing coresets involves sampling data points with probabilities proportional to their so-called sensitivity scores, which quantify the worst-case impact of a point on a given attribute of interest (Phillips, 2017). Take a classification task as an example, this attribute is often a smooth convex approximation of the misclassification rate (Huang et al., 2024). However, the ultimate objective in wide-area scene understanding extends beyond only optimizing the loss function on observed training data to understanding classifier performance on unseen data. This leads to two limitations of existing coreset selection strategies in wide-area scene understanding: (i) their effect-only selection approach, constrained by predetermined partitions, may yield regions that are neither sufficiently small nor dense, thus impeding optimization; and (ii) they tend to overlook the interrelationships between regions, hindering the accurate prediction.
@@ 4.2. Hierarchical Coresets Selection
Specifically, we first uniformly divide the input image $X\in\mathbb{R}^{w\times h\times c}$ into $N_{p}\times N_{p}$ large-scale patches (set as $4\times 4$ in this work), with each patch considered as a candidate region, denoted as =$\mathcal{R}=\{r_{ij}\},i,j=1,\dots,N$. Next, we compute a score $S(r_{ij})$ for each region $r_{ij}$ based on a predefined importance function $S(\cdot)$ (see Eq.7). This function comprehensively integrates factors such as utility, representativeness, robustness, and synergy effects. We then rank all candidate regions by their scores and select the top $K$ regions to form an initial candidate set $\mathcal{R}_{s}$. During this selection, we simultaneously verify that each chosen region satisfies the core set definition, ensuring the resulting core set closely approximates the loss of the entire image, i.e., ensuring $|\mathcal{L}(X)-\mathcal{L}(X_{s})|\leq\epsilon\,\mathcal{L}(X)$ ($\epsilon=0.98$ in this work). After identifying the preliminary candidate set, we further subdivide each selected region $r\in\mathcal{R}_{s}$ into patches and repeat the evaluation and selection process described above. Notably, since we have introduced a synergistic scoring component $S_{sy}$, each chosen region’s combined performance with adjacent patches is explicitly assessed. This strategy facilitates boundary refinement and ensures that the final selected regions collectively meet the core set requirements. Finally, we aggregate the finely screened candidate regions obtained within each coarsely selected patch to construct the final core set $X_{s}=\bigcup_{r\in\mathcal{R}_{s}}r$. Based on this, HCS achieves an efficient selection process that progressively refines from the full image to local regions, from coarse to fine, obtaining a precise coreset for accurate wide-area scene understanding.
@@ 5. Theoretical Analysis
[…]
@@ 6.4. Ablation Study
Impact of Selection Mechanism. HCS employs a hierarchical search to ensure the coreset accuracy while reducing the computational cost (Subsection 4.2). To evaluate its effectiveness, we conduct an ablation experiment, i.e., compare the selection strategy with pixel-wise search. The pixel-wise search is considered to refine the smallest and accurate regions, but it may cause huge computational overhead. Using CLIP and taking the cost of HCS as the benchmark (1×1$\times$1 ×), we record the performance of both methods on NWPU-RESISC45. Figure 4 shows that the computational cost of HCS is less than half that of the pixel-wise search, while the accuracy is nearly identical, demonstrating its superiority.
@@ Acknowledgements.
=$\mathcal{R}=\{r_{ij}\}$ The set of regions $r_{ij}$, $i,j=1,\dots,N$
@@ Appendix A Proofs
[…]
@@
[…]
@@
[…]
@@
(13) =$\mathcal{R}=\Big{\{}\{x\in\mathcal{X}:A(\theta;x)>t\}:\,\theta\in\Theta,\;t\geq 0\Big{\}}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ E.1. Comparison with Other Selection Methods
ViT-B+HCS 17.2 1×1$\times$1 ×
ViT-B+DAUS (Wang et al., 2024a) 13.1 3.2×3.2$\times$3.2 ×
ViT-B+MMTS (Ataee Tarzanagh et al., 2023) 11.9 2.9×2.9$\times$2.9 ×
ViT-B+SCS (Har-Peled and Mazumdar, 2004) 8.6 1.7×1.7$\times$1.7 ×
ViT-B+ESC (Chai et al., 2023) 12.5 2.1×2.1$\times$2.1 ×
ViT-B+al-seg (Vepa et al., 2025) 9.2 2.4×2.4$\times$2.4 ×
CLIP+HCS 21.4 1×1$\times$1 ×
CLIP+DAUS (Wang et al., 2024a) 17.3 3.8×3.8$\times$3.8 ×
CLIP+MMTS (Ataee Tarzanagh et al., 2023) 15.2 3.4×3.4$\times$3.4 ×
CLIP+SCS (Har-Peled and Mazumdar, 2004) 9.1 1.6×1.6$\times$1.6 ×
CLIP+ESC (Chai et al., 2023) 15.0 1.7×1.7$\times$1.7 ×
CLIP+al-seg (Vepa et al., 2025) 13.4 2.7×2.7$\times$2.7 ×
@@ E.2. Model Efficiency and Trade-off Performance
Since the proposed HCS is a plug-and-play method, in order to ensure its practicability, we explore the balance between model performance and efficiency. We compare the trade-off performance of multiple baselines before and after using our HCS. We select TrashCan as the benchmark dataset and the semantic segmentation models as the baselines. It is worth noting that since HCS improves the model’s scene understanding performance at test time without requiring fine-tuning, we do not use training time as an evaluation metric. Instead, we assess the per-image inference speed during deployment, using the embedded baseline model as the reference (i.e., 1×1$\times$1 ×). The results illustrated in Figure 6 show that introducing HCS achieves great performance with acceptable computational cost (less than 1.3$\times$).

### v2

@@ Abstract.
Scene understanding is one of the core tasks in computer vision, aiming to extract semantic information from images to identify objects, scene categories, and their interrelationships. Although advancements in Vision-Language Models (VLMs) have driven progress in this field, existing VLMs still face challenges in adaptation to unseen complex wide-area scenes. To address the challenges, this paper proposes a Hierarchical Coresets Selection (HCS) mechanism to advance the adaptation of VLMs in complex wide-area scene understanding. It progressively refines the selected regions based on the proposed theoretically guaranteed importance function, which considers utility, representativeness, robustness, and synergy. Without requiring additional fine-tuning, HCS enables VLMs to achieve rapid understandings of unseen scenes at any scale using minimal interpretable regions while mitigating insufficient feature density. HCS is a plug-and-play method that is compatible with any VLM. Experiments demonstrate that HCS achieves superior performance and universality in various tasks. The code is available at https://wangjingyao07.github.io/HCS.github.io/.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3754707submissionid: 204ccs: Computing methodologies Scene understandingccs: Computing methodologies Transfer learningccs: Computing methodologies Feature selection
@@ 1. Introduction
Despite their robust representational capabilities (Fu et al., 2024; Cao et al., 2024; Wang et al., 2023b), existing VLMs struggle with wide-area scenes. These scenes, spanning expansive environments such as deep-sea regions, remote sensing imagery, and areas with complex geographical features (Greene and Oliva, 2009; Gu et al., 2019), differ substantially from the urban or indoor settings typical of VLM pre-training (Naseer et al., 2018; Grant and Flynn, 2017). The challenges primarily arise from two interrelated factors. First, the extensive coverage inherent to wide-area scenes introduces an unprecedented diversity of object types (e.g., unknown species of sea snakes, ascidians, or sea urchins in oceans) that are rarely encountered in standard datasets. Second, although with various species, the data may exhibit a pronounced long-tail and sparse distribution in some areas. For instance, while sedimentary rocks might dominate the seabed, rarer features such as sponges appear only sporadically, leading to an overrepresentation of homogenized, high-frequency features. One might argue that existing VLMs boast an abundance of parameters and deeper network architectures that facilitate the capture of intricate and global semantic information. However, these advantages do not inherently overcome the data challenges present in wide-area scenes. In fact, the models may overly focus on globally uniform, high-frequency features. Recent studies have proposed sample partitioning and multi-scale attention mechanisms (Shi et al., 2024; Wang, 2024; Zhang et al., 2024a; Wang et al., 2024c) to address these shortcomings; however, their effectiveness is heavily dependent on the quality of regional segmentation and scale priors, and they often neglect the relationships between different regions. Furthermore, the substantial computational resources required to process the massive volumes of visual data further restrict the practical deployment of such models. Therefore, while pre-training endows VLMs with impressive capabilities, the unique challenges of wide-area scenes pose significant obstacles to their effective adaptation.
@@
To tackle this, we aim to propose a method that can adaptively refine the relationships between scene regions and model decisions in a coarse-to-fine manner for complex wide-area scene understanding. This allows the VLMs to quickly grasp scenes of any scale or complexity using only a few interrelated and interpretable regions. We begin by investigating the interpretable regional characteristics that support effective selection in wide-area scenes. Inspired by data compression theory (Storer, 1987; Sayood, 2017), we propose that the essence of feature selection is to find a small, weighted subset of the original dataset, ensuring that learning tasks on this subset yields an approximately optimal solution. Accordingly, accurate scene understanding requires that the model establish appropriate evaluation criteria tailored to specific tasks, and filter out regions that critically influence decision-making to achieve the key subset, i.e., interpretable regions. To achieve this, we incorporate coreset theory (Har-Peled and Mazumdar, 2004), which enables the construction of a compact subset of input data, i.e., coreset, that serves as a proxy for the whole dataset while maintaining minimal deviation in accuracy. By reducing the data volume, coresets enhance scalability and efficiency. A common approach for constructing coresets involves sampling data points with probabilities proportional to their so-called sensitivity scores, which quantify the worst-case impact of a point on a given attribute of interest (Phillips, 2017). Take a classification task as an example, this attribute is often a smooth convex approximation of the misclassification rate (Huang et al., 2024). However, the ultimate objective in wide-area scene understanding extends beyond only optimizing the loss function on observed training data to understanding classifier performance on unseen data. This leads to two limitations of existing coreset selection strategies: (i) their effect-only selection, bounded by fixed partitions, often produces regions that are too large or sparse, impairing optimization; and (ii) they typically ignore inter-region relationships, reducing prediction accuracy.
@@ 4.2. Hierarchical Coresets Selection
Specifically, we first uniformly divide the input image $X\in\mathbb{R}^{w\times h\times c}$ into $N_{p}\times N_{p}$ large-scale patches (set as $4\times 4$ in this work), with each patch considered as a candidate region, denoted as $\mathcal{R}=\{r_{ij}\},i,j=1,\dots,N$. Next, we compute a score $S(r_{ij})$ for each region $r_{ij}$ based on a predefined importance function $S(\cdot)$ (see Eq.7). This function comprehensively integrates factors such as utility, representativeness, robustness, and synergy effects. We then rank all candidate regions by their scores and select the top $K$ regions to form an initial candidate set $\mathcal{R}_{s}$. During this selection, we simultaneously verify that each chosen region satisfies the core set definition, ensuring the resulting core set closely approximates the loss of the entire image, i.e., ensuring $|\mathcal{L}(X)-\mathcal{L}(X_{s})|\leq\epsilon\,\mathcal{L}(X)$ ($\epsilon=0.98$ in this work). After identifying the preliminary candidate set, we further subdivide each selected region $r\in\mathcal{R}_{s}$ into patches and repeat the evaluation and selection process described above. Notably, since we have introduced a synergistic scoring component $S_{sy}$, each chosen region’s combined performance with adjacent patches is explicitly assessed. This strategy facilitates boundary refinement and ensures that the final selected regions collectively meet the core set requirements. Finally, we aggregate the finely screened candidate regions obtained within each coarsely selected patch to construct the final core set $X_{s}=\bigcup_{r\in\mathcal{R}_{s}}r$. Based on this, HCS achieves an efficient selection process that progressively refines from the full image to local regions, from coarse to fine, obtaining a precise coreset for accurate wide-area scene understanding.
@@ 5. Theoretical Analysis
Let $X\in\mathbb{R}^{w\times h\times c}$ be a wide-area scene image and $\mathcal{X}$ be the set of minimal units extracted from $X$. Let $P$ be a probability measure on $\mathcal{X}$ and define the loss of a fixed model $f_{\theta}$ on $X$ as Eq.1. Assume there exists an upper importance function $s:\mathcal{X}\to(1,\infty)$ for $\ell(\cdot,\theta)$ with total score $S=\int_{\mathcal{X}}s(x)\,dP(x)$ for $X_{s}$. Through HCS, if we select $X_{s}\subset\mathcal{X}$ that satisfy $m\geq g(\epsilon,S,\delta)=\frac{2S}{\epsilon^{2}}\Bigl(8C^{2}+S\ln\frac{2}{\delta}\Bigr)$ and assign weights $\nu(x)=\frac{S}{m\,s(x)}$, then with probability at least $1-\delta$ it holds for all $\theta$ achieve $\left|\mathcal{L}(\theta)-\sum_{x\in X_{s}}\nu(x)\ell(x,\theta)\right|\leq\epsilon\,\mathcal{L}(\theta)$.
@@ 6.4. Ablation Study
Impact of Selection Mechanism. HCS employs a hierarchical search to ensure the coreset accuracy while reducing the computational cost (Subsection 4.2). To evaluate its effectiveness, we conduct an ablation experiment, i.e., compare the selection strategy with pixel-wise search. The pixel-wise search is considered to refine the smallest and accurate regions, but it may cause huge computational overhead. Using CLIP and taking the cost of HCS as the benchmark ($1\times$), we record the performance of both methods on NWPU-RESISC45. Figure 4 shows that the computational cost of HCS is less than half that of the pixel-wise search, while the accuracy is nearly identical, demonstrating its superiority.
@@ Acknowledgements.
$\mathcal{R}=\{r_{ij}\}$ The set of regions $r_{ij}$, $i,j=1,\dots,N$
@@ Appendix A Proofs
Proof. Recall that our setting is as follows. Let $X\in\mathbb{R}^{w\times h\times c}$ be a wide-area scene image and let $\mathcal{X}$ denote the set of minimal units extracted from $X$. The full loss of a fixed model $f_{\theta}$ is defined by $\mathcal{L}(\theta)=\int_{\mathcal{X}}\ell(x,\theta)\,dP(x)$ where $P$ is a probability measure on $\mathcal{X}$. We assume there exists an upper importance function $s:\mathcal{X}\to(1,\infty)$ for $\ell(\cdot,\theta)$ with total score $S=\int_{\mathcal{X}}s(x)\,dP(x)$ via our HCS procedure, we obtain a coreset $X_{s}\subset\mathcal{X}$ of size $m$ and assign each selected unit the weight $\nu(x)=\frac{S}{m\,s(x)}$. Our goal is to show that, with probability at least $1-\delta$, for all $\theta$ the weighted loss estimate $\hat{\mathcal{L}}(\theta)=\sum_{x\in X_{s}}\nu(x)\,\ell(x,\theta)$ satisfies $\Big|\mathcal{L}(\theta)-\hat{\mathcal{L}}(\theta)\Big|\leq\epsilon\,\mathcal{L}(\theta)$.
@@
(10) $\int_{\mathcal{X}}A(\theta;x)\,dP(x)=\frac{S}{\mathcal{L}(\theta)}\int_{\mathcal{X}}\frac{\ell(x,\theta)}{s(x)}\,dP(x).$
@@
(11) $\begin{split}\int_{\mathcal{X}}A(\theta;x)\,dQ(x)&=\frac{S}{\mathcal{L}(\theta)}\int_{\mathcal{X}}\frac{\ell(x,\theta)}{s(x)}\,dQ(x)\\ &=\frac{1}{\mathcal{L}(\theta)}\int_{\mathcal{X}}\ell(x,\theta)\,dP(x)=1.\end{split}$
@@
(13) $\mathcal{R}=\Big\{\{x\in\mathcal{X}:A(\theta;x)>t\}:\,\theta\in\Theta,\;t\geq 0\Big\}$
@@
(15) $\Biggl|\int_{\mathcal{X}}A(\theta;x)\,dQ(x)-\int_{\mathcal{X}}A(\theta;x)\,d\nu(x)\Biggr|\leq\epsilon\Biggl(1+S\cdot\frac{1}{S}\Biggr)=2\epsilon.$
@@
(17) $\varphi_{\lambda,h}(x,y)=\min_{\hat{y}\in\mathcal{Y}}\Bigl\{h(x,y)-h(x,\hat{y})+\lambda\,\mathbf{1}_{\{\hat{y}=y\}}\Bigr\},$
@@
(30) $\varphi_{\lambda,h}(x_{i},y_{i})=h(x_{i},y_{i})-\max_{z\in\mathcal{Y}}\Bigl\{h(x_{i},z)-\lambda\,\mathbf{1}_{\{z=y_{i}\}}\Bigr\}.$
@@
$=\frac{1}{m}\mathbb{E}_{S,\sigma}\left[\sup_{h\in\mathcal{H}}\sum_{i=1}^{m}\sigma_{i}\Bigl(h(x_{i},y_{i})-\max_{z\in\mathcal{Y}}\{h(x_{i},z)-\lambda\,\mathbf{1}_{\{z=y_{i}\}}\}\Bigr)\right]$
@@
$\quad+\frac{1}{m}\mathbb{E}_{S,\sigma}\left[\sup_{h\in\mathcal{H}}\sum_{i=1}^{m}\sigma_{i}\,\max_{z\in\mathcal{Y}}\{h(x_{i},z)-\lambda\,\mathbf{1}_{\{z=y_{i}\}}\}\right].$
For the first term, we express $h(x_{i},y_{i})=\sum_{z\in\mathcal{Y}}h(x_{i},z)\,\mathbf{1}_{\{z=y_{i}\}}$ and then apply linearity and symmetry of the Rademacher variables. This yields a bound proportional to a constant (depending on $|\mathcal{Y}|$) times $\mathfrak{R}_{m}\bigl(\Pi_{1}(\mathcal{H})\bigr)$, where $\Pi_{1}(\mathcal{H})$ denotes the projection of $\mathcal{H}$ onto its coordinate functions. A similar bound applies to the second term. Therefore, there exists a constant $C>0$ such that
@@ E.1. Comparison with Other Selection Methods
ViT-B+HCS 17.2 $1\times$
ViT-B+DAUS (Wang et al., 2024a) 13.1 $3.2\times$
ViT-B+MMTS (Ataee Tarzanagh et al., 2023) 11.9 $2.9\times$
ViT-B+SCS (Har-Peled and Mazumdar, 2004) 8.6 $1.7\times$
ViT-B+ESC (Chai et al., 2023) 12.5 $2.1\times$
ViT-B+al-seg (Vepa et al., 2025) 9.2 $2.4\times$
CLIP+HCS 21.4 $1\times$
CLIP+DAUS (Wang et al., 2024a) 17.3 $3.8\times$
CLIP+MMTS (Ataee Tarzanagh et al., 2023) 15.2 $3.4\times$
CLIP+SCS (Har-Peled and Mazumdar, 2004) 9.1 $1.6\times$
CLIP+ESC (Chai et al., 2023) 15.0 $1.7\times$
CLIP+al-seg (Vepa et al., 2025) 13.4 $2.7\times$
@@ E.2. Model Efficiency and Trade-off Performance
[…]